# CP-Agent: Agentic Constraint Programming

### v2

@@ 6.1 Experimental Setup
Our experimental setup executed all 101 problems using CP-Agent, implemented as agentic-python-coder (version 2.2.1), with Claude Sonnet 4.5 (version 20250929). Each problem execution produces a solution file containing the generated CPMpy model and a conversation log recording the complete agent interaction. We validated the solutions by running them against the ground-truth CPMpy models. For satisfaction problems, we verified that the solutions satisfy all ground-truth constraints. For optimization problems, we also compared the objective value to confirm optimality. This approach correctly handles problems with multiple valid solutions while ensuring correctness and accuracy. We ran each experiment three times to account for stochastic variance.
The agentic-python-coder and all experimental data, including problem specifications, ground-truth models, generated solutions, and execution logs, are available online [22]. The data is encrypted (password Ckc1ADsKXRfmzD2) to prevent training data contamination.
@@ 6.2 Results
CP-Agent achieved 100% accuracy on all 101 CP-Bench problems, successfully solving all problems across all three independent runs. All generated solutions were validated against the reference CPMpy models and produced correct and optimal solutions.
Table 1: Tool usage, token consumption, and execution time for all 101 CP-Bench problems (averaged over three independent runs). Column ex shows the number of python_exec calls (iterative refinement cycles). Columns in/out show input/output tokens in thousands. Column time shows execution time in seconds (rounded).
@@
Token consumption averaged 37k input and 3k output tokens per problem. Problem 014 (crossfig) represents an extreme outlier, consuming 737k input tokens due to its complex crossword-figure structure, which requires extensive iterative constraint refinement.
Execution times ranged from 20 seconds to 542 seconds, with an average of 52 seconds per problem across the three runs. The timing variation between the three independent runs (averaging 48, 57, and 52 seconds, respectively) demonstrates the stochastic nature of LLM-based systems.
Table 1 summarizes computational cost; solution quality is perfect across all runs. The agent called python_exec between 1 and 28 times per problem, with simpler problems requiring fewer iterations and complex scheduling or optimization problems requiring more exploration. Problem 014 (crossfig) required the most Python executions (28 average), followed by 009 (basketball) with 24 executions, reflecting their combinatorial complexity.
To illustrate how the agent approaches constraint modeling, we examine two representative runs that demonstrate different problem-solving patterns.
@@ 6.3 Problem 061: Cryptarithmetic Puzzle
The cryptarithmetic problem required assigning unique digits (0–9) to letters in a 20-digit addition equation, ensuring the arithmetic holds while each letter maps to a distinct digit. The agent’s solution process demonstrates strategic adaptation when an initial approach proves unworkable.
The first four executions built the foundation: defining ten integer variables (one per letter), adding an AllDifferent constraint, mapping variables to their positions in the three numbers, and constraining leading digits to be non-zero. This setup proceeded smoothly, establishing the basic problem structure.
The fifth execution attempted to model the problem as a single unit by treating each 20-digit number as a single large integer variable. The agent constructed expressions like num1 = letter[0]*10**19 + letter[1]*10**18 + ... for all three numbers and added a constraint num1 + num2 == result. When the model tried to solve, the solver raised a TypeError: __init__(): incompatible constructor arguments because variable domains with bounds near $10^{20}$ exceeded the solver’s integer representation capacity.
Recognizing the fundamental incompatibility between this modeling approach and solver limitations, the agent abandoned the large-integer strategy entirely. The agent redesigned the model, introducing binary carry variables and modeling addition column by column, thereby mimicking manual arithmetic:
@@
This column-wise formulation ensured that all intermediate values remained within standard integer bounds, enabling the solver to determine the correct digit assignment.
@@ 6.4 Problem 082: Wolf, Goat, and Cabbage
While the above example demonstrates error recovery, this river-crossing puzzle illustrates the agent’s behavior when proceeding without failures. The agent built the temporal model through four distinct phases.
The first phase established state variables: wolf_pos[t], goat_pos[t], cabbage_pos[t], and boat_pos[t], which track locations (shore 0 or 1) across an 8-step timeline. The second phase fixed boundary conditions—all entities start at shore 0 and end at shore 1. The third phase encoded the movement mechanics: the boat crosses at each step, the farmer transports at most one item per trip, and item movements coincide with boat movements. The fourth phase added safety constraints:
@@
These constraints forbid the goat and cabbage (or wolf and goat) from sharing a shore without the farmer present. The systematic layering—state, boundaries, movement, and safety—reflects a deliberate construction process where the agent tests each phase before proceeding to the next. The model solved correctly on the first execution.
These two problems reveal complementary behaviors: complete redesigns when facing fundamental obstacles (Cryptarithmetic) and methodical incremental construction when the path is clear (Wolf, Goat, and Cabbage). The agent adapts its approach rather than following a fixed procedure.
@@ 7 Model Generalization
To evaluate whether the approach generalizes beyond Claude Sonnet 4.5, we tested six frontier models on ten representative problems with three runs each. The models span four providers: Anthropic (Sonnet 4.5, Opus 4.5, Haiku 4.5), OpenAI (GPT-5.1-codex-mini), xAI (Grok-code-fast-1), and DeepSeek (v3.2). We excluded Google’s Gemini models due to infrastructure issues during the experimental period. All models used a temperature of 0.0 with a 20-minute timeout per problem.
Table 2 shows the results. All tested models achieved a success rate of 80–100% using this architecture. Sonnet 4.5 maintained perfect accuracy, while Opus 4.5 and GPT-5.1 each reached 97%. Haiku 4.5, Grok, and DeepSeek achieved 80–87% accuracy, with most failures occurring on Problem 009 (basketball), although sporadic failures also occurred on other problems. Basketball requires coordinating nine constraint categories, including mirroring schemes, home/away patterns, rival pairings, and date-specific requirements—the highest constraint density in this subset.
Some runs hit the 20-minute timeout before completion. Review of execution logs confirms these agents were not converging toward correct solutions, suggesting the timeout did not cut short otherwise successful attempts.
These results show the architecture generalizes across model families. Success correlates with model capability rather than provider: the agentic workflow provides value at all capability levels tested. Prompt optimization frameworks, such as DSPy [10], could automatically tune the project prompt for specific models, potentially narrowing the performance gap observed with other models.
@@ 8 Task Management Tool
Explicit task tracking for LLM agents—maintaining a mutable list of pending, in-progress, and completed tasks—was introduced by BabyAGI [15] and formalized in plan-and-solve prompting [27]. The todo_write tool in our coder implements this pattern. Our architecture requires only todo_write, not a corresponding todo_read: without history compaction, all previous tool calls remain in the conversation context, so the agent can always see its earlier task updates. To evaluate whether explicit task management benefits constraint modeling, we conducted an ablation study comparing performance with and without this tool.
We used a smaller model (Claude Haiku 4.5) than our default model (Claude Sonnet 4.5) to reveal potential differences. We selected 40 problems, stratified by difficulty: 20 easy problems (solved in 27 seconds or less by Sonnet 4.5) and 20 hard problems (requiring 58 seconds or more). For each problem, we executed three independent runs with todo_write disabled and three with it enabled. Table 3 presents the results.
@@
On easy problems, todo_write increased token consumption by 191% (from 15K to 44K average) and execution time by 111% (from 18s to 38s), while improving success by just one run (59/60 to 60/60). On hard problems, todo_write reduced success from 52/60 to 45/60 while still adding overhead: 41% more tokens (209K to 295K) and 27% longer execution (98s to 124s).
Examining individual problems reveals the pattern. The task management tool helped three previously failing hard problems—crossfig, sports, and progressive—by providing structure for extended problem-solving. However, it broke three problems that worked without it: basketball, crew, and wolf_goat all dropped from 3/3 to 0/3. The task-tracking overhead appears to distract the agent from solutions it would otherwise find.
9 Discussion
Our experiments yield several key observations regarding the design and evaluation of agentic systems for constraint modeling.
The ablation study, which shows that a short prompt outperforms an 800-line version, suggests that for foundation models with constraint programming knowledge in their training data, the prompt’s role is to set high-level strategy and details on output format rather than provide exhaustive procedural instructions. Prompt compression research supports this finding: shorter, denser prompts can match or exceed verbose prompts [9]. Excessive detail may limit the agent’s ability to solve problems effectively.
The primary evaluation used Claude Sonnet 4.5. Multi-model testing on a representative 10-problem subset confirms that the architecture generalizes: six models from four providers achieved success rates of 80–100%, with performance correlating with model capability.
This pattern extends to built-in planning mechanisms. The optional todo_write tool, intended to help the agent manage tasks, reduced overall success when tested with a smaller model. The effect was asymmetric: todo_write helped on already-failing problems but broke problems that worked without it. For problems where the entire solution fits in a single script, the overhead of maintaining an explicit task list is more distracting than beneficial. When state is fully captured in code and execution output, the agent’s internal reasoning seems to suffice without externalized planning.
@@ 10 MCP Server
The Model Context Protocol (MCP) provides a standardized interface for connecting language models with external tools and data sources [8]. Introduced by Anthropic in late 2024 [1], MCP has become a de facto standard for agent-tool interaction, with adoption by major AI platforms including OpenAI and integration into development environments such as Cursor, Claude Code, VS Code (via Cline), Windsurf, and OpenAI Codex CLI. In December 2025, MCP governance transferred to the Agentic AI Foundation under the Linux Foundation.
While our standalone CP-Agent excels at batch processing, interactive development benefits from human guidance through complex modeling decisions. We expose the same persistent IPython kernel as an MCP server (ipython_mcp) for integration with Model Context Protocol-compatible clients such as Claude Desktop or Cursor. The server provides python_exec for code execution, the same tool central to our standalone agent, along with three tools for controlling and monitoring the coding infrastructure: python_reset to clear state and install packages to the virtual environment, python_status to query session information, and python_interrupt to halt long-running computations. This interactive mode also serves an educational purpose: users can observe how natural language constraints are formalized into CPMpy models, request explanations, and refine models based on solver feedback. Once developed, the generated code can be saved for deployment in other contexts.
Our MCP integration complements the system MCP-Solver [21], which provides specialized tools for constraint modeling (MiniZinc, PySAT, Z3, ASP). In contrast, the ipython_mcp server offers general-purpose Python execution, where domain adaptation occurs through prompts rather than tool specialization.
@@ 11 Conclusion
CP-Agent solves all 101 problems in the clarified CP-Bench benchmark, demonstrating that agentic workflows with execution feedback can reliably automate constraint modeling from natural language.
The agent architecture combines the ReAct framework with a persistent IPython kernel. Our ablation study shows that a minimal project prompt is sufficient, and that explicit task management tools generally add overhead for focused modeling tasks. However, they can help maintain focus during longer problem-solving sessions. For foundation models with relevant domain knowledge, concise guidance and stateful execution prove more effective than detailed procedural scaffolding or explicit planning mechanisms.
The architecture’s separation of domain expertise (prompts) from execution infrastructure (the agent) allows adaptation to other scientific computing domains. Follow-up work on ASP-Bench [23] demonstrated excellent performance on 128 answer set programming problems, utilizing the identical architecture with only a domain-specific project prompt, thereby confirming that our approach generalizes across declarative programming paradigms.
Acknowledgments
This research was supported by the Austrian Science Fund (FWF) projects 10.55776/P36420 and 10.55776/COE12.

### v3

@@ 6.1 Experimental Setup
We ran all our experiments on all 101 CP-Bench problems three times, with CP-Agent implemented as agentic-python-coder (version 2.2.1) using Claude Sonnet 4.5 (version 20250929) as the language model for our main experiment. The models used in the model generalization and the task management experiments are described below. For each problem, the agent produced a solution file containing the generated CPMpy model, and we recorded a conversation log with the complete agent interaction. We tested all solutions against the ground-truth CPMpy models. For each satisfaction problem, we tested whether the solution satisfies all ground-truth constraints. For optimization problems, we compared the objective value of the generated solution with an optimal pre-computed ground-truth value to confirm optimality. Note that a problem may have several correct and optimal solutions, hence the ground-truth model must not contain symmetry-breaking constraints. We always report results averaged across three runs to account for randomness.
The agentic-python-coder, the problem specifications and ground-truth models to reproduce our experiments, and the generated solutions and execution logs, are available online [22]. The data is encrypted (password Ckc1ADsKXRfmzD2) to avoid training data contamination.
@@ 6.2 Results
Our experimental results show that CP-Agent achieves 100% accuracy, successfully solving all 101 CP-Bench problems across all three independent runs. We validated all generated solutions against the ground-truth models and confirmed that all solutions were correct and optimal.
Our results are summarized in the table on the next page: tool usage, token consumption, and execution time for all 101 CP-Bench problems (averaged over three independent runs). Column ex shows the number of python_exec calls (iterative refinement cycles). Columns in/out show input/output tokens in thousands. Column time shows execution time in seconds (rounded).
The number of python_exec calls varies and depends on the structure of the problem, ranging from 1 to 28, with simpler problems requiring fewer iterations and complex scheduling or optimization problems requiring more exploration. Problem 014 (crossfig) stands out with an unusually large number of Python executions (28 average), followed by 009 (basketball) with 24 executions, reflecting their combinatorial complexity.
@@
The table also shows the token consumption. On average we have 37k input and 3k output tokens per problem. As expected, Problem 014 (crossfig), the problem with the largest number of exec calls has also the largest token consumption (737k input tokens).
The specifics of execution times differ from problem to problem, ranging from 20 seconds to 542 seconds, with an average of 52 seconds across the three runs. The average run time per problem in each of the three runs is 48, 57, and 52 seconds, respectively, which demonstrates the stochastic nature of LLM-based systems, but can also reflect variety in API response time and network speed.
We present representative runs on two problems to illustrate how the agent approaches constraint modeling, with different problem-solving patterns.
@@ 6.3 Problem 061: Cryptarithmetic Puzzle
The problem is to assign letters with unique digits (0–9) such that the given 20-digit addition equation holds. The log reveals how the agent adapts when an initial approach proves unworkable.
In the first four executions, the agent introduces integer variables (one per letter), an AllDifferent constraint, mappings from variables to their positions in the three numbers, and constraints forcing leading digits to be non-zero. This completes the description of the problem structure.
The fifth execution attempts to treat each 20-digit number as a single large integer variable. The agent constructs expressions like num1 = letter[0]*10**19 + letter[1]*10**18 + ... for all three numbers and adds the constraint num1 + num2 == result. This approach fails because variable domains of size $10^{20}$ exceed the solver’s integer representation capacity, causing a TypeError: __init__(): incompatible constructor arguments.
With this feedback, the agent redesigns the model, introducing binary carry variables and models addition column by column, thereby mimicking manual arithmetic:
@@
This way, all intermediate values are kept within standard integer bounds, enabling the solver to determine the correct digit assignment.
@@ 6.4 Problem 082: Wolf, Goat, and Cabbage
The above example demonstrated error recovery; this well-known river-crossing puzzle illustrates the agent’s behavior when proceeding without failures.
The first step is to establish state variables: wolf_pos[t], goat_pos[t], cabbage_pos[t], and boat_pos[t], which track locations (shore 0 or 1) across an 8-step timeline. The second step of the construction fixes boundary conditions—all entities start at shore 0 and end at shore 1. The third step encodes the movement mechanics: the boat crosses at each step, the farmer transports at most one item per trip, and item movements coincide with boat movements. The fourth step adds safety constraints:
@@
These constraints ensure that the goat and cabbage (or wolf and goat) do not share a shore without the farmer present. The systematic layering—state, boundaries, movement, and safety—naturally reflects the construction process. Each phase is traversed in order to verify correctness before proceeding to the next. The execution completes successfully on the first attempt.
These two behaviors of the agent can be observed in different phases of problem solving: complete redesigns when facing fundamental obstacles and methodical incremental construction when the path is clear. The agent uses adaptive approaches rather than fixed workflows.
@@ 7 Model Generalization
To study whether the approach generalizes beyond Claude Sonnet 4.5, we evaluated six frontier models on ten representative problems, with three runs per model. The models come from four model makers: Anthropic (Sonnet 4.5, Opus 4.5, Haiku 4.5), OpenAI (GPT-5.1-codex-mini), xAI (Grok-code-fast-1), and DeepSeek (v3.2). Due to infrastructure issues during the experimental period, we did not test with Google’s Gemini models. We ran each problem with a 20-minute timeout, using a temperature of 0.0 for each model.
We give the results in Table 1. The tested models achieved a success rate of 80–100%. Sonnet 4.5 maintained perfect accuracy. Opus 4.5 and GPT-5.1 both reached 97%. Haiku 4.5, Grok, and DeepSeek achieved 80–87% accuracy; most failures occurred on Problem 009 (basketball), though sporadic failures also affected other problems. Basketball requires coordinating nine constraint categories, including mirroring schemes, home/away patterns, rival pairings, and date-specific requirements—the highest constraint density in this subset.
For some runs, the solver timed out after 20 minutes. Looking at execution logs, we see that these agents were not converging toward correct solutions, suggesting the timeout did not cut short otherwise successful attempts.
Our results show that the architecture generalizes across model families. This indicates that our agentic workflow generalizes from our default model, Sonnet. Opus, considered superior to Sonnet, showed slightly weaker performance, possibly because the prompt was optimized for Sonnet. Prompt optimization frameworks such as DSPy [10] could be used to automatically tune the project prompt for specific models, potentially narrowing the performance gap observed with other models.
@@ 8 Task Management Tool
Explicit task tracking for LLM agents—maintaining a mutable list of pending, in-progress, and completed tasks—was introduced by BabyAGI [15] and formalized in plan-and-solve prompting [27]. Our implementation provides a tool that realizes this pattern. Our architecture requires only todo_write, not a corresponding todo_read: without history compaction, all previous tool calls remain in the conversation context, so the agent can always see its earlier task updates. To evaluate whether explicit task management benefits constraint modeling, we compare our method with and without this tool.
For this experiment, we use a smaller model (Claude Haiku 4.5) than our default (Claude Sonnet 4.5) to reveal potential differences. We consider the following 40 problems, stratified by difficulty: 20 easy problems (solved in 27 seconds or less by Sonnet 4.5) and 20 hard problems (requiring 58 seconds or more). We run each problem three times with todo_write disabled and three times with it enabled. The results are summarized in Table 2.
@@
On easy problems, compared with the baseline, todo_write results in a significant increase in token consumption (191%, from 15K to 44K average) and execution time (111%, from 18s to 38s), while improving success by just one run. On hard problems, todo_write even reduced success from 52/60 to 45/60 while still adding overhead in terms of token usage and longer execution time.
A closer look at individual problems reveals a pattern. The task management tool provides structure for extended problem-solving, which helps three previously failing hard problems—crossfig, sports, and progressive. In contrast, it causes failures on three problems that worked without it. The task-tracking overhead appears to distract the agent from finding solutions in some cases.
@@ 9 MCP Server
The Model Context Protocol (MCP) is a standardized interface for connecting language models with external tools and data sources [8]. MCP was initially introduced by Anthropic in late 2024 [1] as a tool for agent-tool interaction, and has since been adopted by major AI platforms and integrated into development environments. In December 2025, MCP governance transferred to the Linux Foundation.
To make our framework accessible from other agentic systems, beyond our own ReAct agent, we define the same persistent IPython kernel system as an MCP server (ipython_mcp). This allows it to be integrated with any MCP-compatible client. The MCP server is equipped with python_exec for code execution, the same tool central to our standalone agent. It also comprises three tools for controlling the coding infrastructure: python_reset to clear state and install packages, python_status to query session information, and python_interrupt to halt long-running computations. This interactive mode also serves an educational purpose: users can observe how natural language constraints are formalized into CPMpy models, request explanations, and refine models based on solver feedback. Once developed, the generated code can be deployed to other contexts.
This approach complements the MCP-Solver [21], which provides specialized tools for constraint modeling (MiniZinc, PySAT, Z3, ASP). In contrast to the MCP-Solver, our new ipython_mcp server provides general-purpose Python execution, with domain adaptation driven by prompts rather than tool specialization.
@@ 10 Conclusion
Our experiments provide several insights into the design and evaluation of agentic systems for constraint modeling.
Our results show that agentic workflows with execution feedback can reliably automate constraint modeling from natural language.
Our main experiment demonstrates that a short prompt performs very well (even outperforms an 800-line earlier version). This suggests that, for foundation models with constraint-programming knowledge, the role of the prompt is to set high-level strategy and the output format rather than to provide exhaustive procedural instructions. This is in line with recent results on prompt compression showing that shorter, denser prompts can match or exceed verbose prompts [9]. In terms of procedural detail, excessive instruction might reduce the agent’s ability to solve problems effectively.
Our model generalization experiments show that our agentic architecture is robust: six models from four providers achieved success rates of 80–100%, with performance correlating with model capability.
Our experiment with the task management tool gives a mixed picture. While the tool reduces the overall success rate and significantly increases time and token usage, it was helpful for a few problems. Based on this outcome, we turn off the tool by default but keep it in the codebase for further testing, as there might be problem families where it shows a positive overall performance.
Finally, we reported on further integration possibilities of our system via the Model Context Protocol.
We believe that separating domain expertise from execution infrastructure, as realized in our system, can be beneficial in other scientific computing domains. In follow-up work, we used our system to transform natural language descriptions into logic programs [23], using the same architecture with only a domain-specific project prompt.